can i please get help with this problem
A survey found that one out of four people say that they have visited a doctor in any given month. If 14 people are selected at random, find the probability that exactly 6 will have visited a doctor last month. Please round your answer to three decimal places

Answer by stanbon(75887)

A survey found that one out of four people say that they have visited a doctor in any given month. If 14 people are selected at random, find the probability that exactly 6 will have visited a doctor last month. Please round your answer to three decimal places
-----
Binomial Problem with n = 14 and p = 1/4
---
P(x=6) = 14C6(1/4)^6(3/4)^8 = binompdf(14,1/4,6) = 0.073
============
Cheers,
Stan H.
=============